HUBBLE SPACE TELESCOPE DAILY REPORT #2351 PERIOD COVERED: 0000Z (UTC) 04/14/99 - 0000Z (UTC) 04/15/99 Daily Status Report as of 105/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7949 (CCD Dark and Bias Monitor -- Cont. to June 99) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a CCD dark calibration and a bias monitoring program. It was also scheduled to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7790 (WF/PC-2 Narrow-Band Imagery of the Carina Nebula {NGC-3372}) The WF/PC-2 was used to make narrow band imagery of the Carina Nebula (a.k.a., NGC-3372). The intricate "Keyhole" structure at the center of the Carina Nebula was imaged in several WF/PC-2 narrow-band filters. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 7306 (Halo Microlens Source Systems and their Underlying Stellar Backgrounds) The WF/PC-2 was used to to obtain medium-deep broadband images of source star STAR-0501 for microlensing events, plus one additional star, to determine a} which star was actually lensed, b} the amount of crowding in our ground-based frames, c} if the stars are normal in the CMD, and d} if they lie preferentially at the back of the LMC. The proposal completed with no reported problems. 1.5 Completed Six Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.6 Completed Three Sets of WF/PC-2 6357 (The Nuclear Morphology of Elliptical Galaxies) The WF/PC-2 was used to make images of the early type, large, elliptical galaxies NGC-5308, NGC-4478 and ESO507-27 (taken from a list of 150 galaxies) in order to study their nuclear morphology. This survey will complement Illingworth's snapshot study of later Hubble types. The observations were completed as scheduled, and no anomalies were reported. 1.7 Completed Two Sets of WF/PC-2 6480 (The Butcher-Oemler Effect: Where Have All The Spirals Gone?) The WF/PC-2 was used to make observations of the blue disk-type galaxies A851-Pos6 and A851-Pos7 in order to study the Butcher-Oemler effect. Other observational data show that mergers and tidal interactions were apparently more common in younger clusters; these may contribute to the disappearance of spiral galaxies by the present epoch. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/MA1 7401 (A Critical Test of Crystallization in White Dwarf Stars) The Space Telescope Imaging Spectrograph (MA1) was used to make observations of the white dwarf star BPM37093-2. Theoretical calculations indicate that the massive pulsating DA white dwarf, BPM37093, has a crystallized ion interior. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Operations Notes: The operations team moved to the SAMS backup facility at approximately 104/1700Z in preparation for the first simulations for Servicing Mission 3A (SM3A). This simulation is scheduled to start at 105/1200Z and last until 106/0300Z. Using ROP SR-1, the SSR EDAC error counter was cleared twice. The SSA transmitter was turned on and off per ROP IC-2. Per an operations request, a STIS memory dump of EDAC and EEPROM was executed. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH